solve:
x^2-y^2=4
x+y=1

Answer by stanbon(48568)

1st: x^2-y^2=4
2nd: x+y=1
Divide 1st by 2nd to get:
3rd: x-y=4
-------------
Add 2nd and 3rd to get:
4th: 2x=5
x=5/2
---------
Substitute into 2nd to get:
(5/2)+y=1
y=-3/2
-------------
Solution: x=5/2 ; y=-3/2
---------
Checking the solution:
1st: x^2-y^2=4
(5/2)^2-(-3/2)^2=4
25/4 - 9/4 = 4
16/4 = 4
4=4
-----------
Cheers,
Stan H.